product and role 2 Hound your 12. Using the ABC system, what percentage of the Machining costs is assigned to Product Y and P13. Using the ABC system, what percentage of Machine Setups cost is assigned to Product Y and Product Z? (Round your Percent14. Using the ABC system, what percentage of the Product Design cost is assigned to Product Y and Product Z? Product Y Produc15. Using the ABC system, what percentage of the General Factory cost is assigned to Product Y and Product Z? (Round your PeRequired information The Foundational 15 [LO7-1, LO7-3, L07-4] [The following information applies to the questions displayed
Answer #1
  • Requirement 12 to 15

Product Y

Product Z

Requirement 12

Machine Cost

72.73%

27.27%

Requirement 13

Machine Setup Cost

20.00%

80.00%

Requirement 14

Product Design Cost

50.00%

50.00%

Requirement 15

General Factory Cost

53.03%

46.97%

  • Working

Product Y

Product Z

Requirement 12

Machine Cost

=8000/11000

=3000/11000

Requirement 13

Machine Setup Cost

=60/300

=240/300

Requirement 14

Product Design Cost

=1/2

=1/2

Requirement 15

General Factory Cost

=7000/13200

=6200/13200